A manufacturer is studying the effects of cooking temperature, cooking time, and type of cooking oil for making potato chips. Three different temperatures, 4 different cooking times, and 3 different oils are to be used.
(a) What is the total number of combinations to be studied?
(b) How many combinations will be used for each type of oil?
(c) Discuss why permutations are not an issue in this exercise.
1
Expert's answer
2013-02-12T10:40:20-0500
a) the total numberof combinations to be studied is:
N = Nt*No*Nc
Nt - number of different temperatures
No- number of different cooking times
Nc - number of different oils
N = 3*4*3 = 36
b) number of combinations for each type of oil:
N1o = Nt*Nc
Nt - number of different temperatures
No- number of different cooking times
N1o = 3*4 = 12
c) because the cooking temperature,cooking time, and type ofcooking oil are not dependent, so for us it does not matter in what order
to list them.